OpenStudy (anonymous):

Solve 0.25(7 + 3m) = -0.125 m. -2 2 -40.5 40.5

OpenStudy (anonymous):

it doesn't take that long to reply ..

OpenStudy (kropot72):

First multiply out to remove the brackets giving: 1.75 + 0.75m = -0.125m Next subtract 1.75 from both sides giving: 0.75m = -0.125m - 1.75 Then add 0.125m to both sides giving: 0.875m = -1.75 m = -1.75/0.875 = -2